Find the horizontal and vertical components of a force of 120𝑁 that acts in a direction forming a 58° with the vertical.


1
Expert's answer
2021-04-29T17:52:40-0400

"F = 120 \; N" , "\alpha =58^\circ"

The horizontal component:

"F_x = 120 \cdot \cos 58^\circ = 120 \cdot 0.53 = 63.6\; N"

The vertical component:

"F_y = 120 \cdot \sin 58^\circ = 120 \cdot 0.848 = 101.76 \; N"